if the demand curve id P=500 - 4Q .. what is the value of the marginal revenue when Q =20.
Answer #1

P = 500-48 TR= P8 = 5008 - 40² onk= all - 500-88 28 hehen d=20 MR = 500-8 (20) = 500- 160 = 340 Marginal revenue - 340